OpenStudy (anonymous):

on a certain day in Toronto,Canada the temp. is 15 degrees Celsius (c). USING THE FORMULA F= 9/5 C+32. which temp represents 15 degrees C in degrees Fahrenheit A -9 B 35 C 59 D 85

OpenStudy (anonymous):

What will be your answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do you think answer should be? Hint Put the given value in C

OpenStudy (anonymous):

is it a

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

ugh haha

OpenStudy (anonymous):

F = 9/5(C) + 32 C = 15 F = 9 (15/5) + 32 F = ???

OpenStudy (anonymous):

i got 59 but why did we divide 15 by 5

OpenStudy (anonymous):

because 15 is divisible by 5. If we had, say, 43 then F = 9/5(43) +32

OpenStudy (anonymous):

oh thankyou so much

OpenStudy (anonymous):

Anytime :)